OpenStudy (anonymous):

let f(x)=5x+12. find f^1(x)

OpenStudy (anonymous):

write again clearly

OpenStudy (anonymous):

let f(x)= \[\frac{ 24 }{ 3x-2 } find f(-2)\] @zith

OpenStudy (debbieg):

What did you get, @allie_bear22 ? Plug in -2 for x and simplify. :)

OpenStudy (debbieg):

In the denominator you'll have 3(-2)-2 = ?

OpenStudy (anonymous):

my answer is \[\frac{ 5 }{ x-12 }\]

OpenStudy (anonymous):

no bear put -2 in place of x and then solve

OpenStudy (anonymous):



OpenStudy (debbieg):

I don't think it's clear what the problem is... the function posted in your original question is completely different from the one 2 posts below it, lol.

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

answer is -3

OpenStudy (anonymous):

im sorry i posted it wrong let my try again lol

OpenStudy (anonymous):

let f(x)=5x+12. find \[f^1\](x)

OpenStudy (anonymous):

f'(x)=5

OpenStudy (debbieg):

is it \[\large f ^{-1}(x)\]the inverse function?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

no i think its derivative

OpenStudy (anonymous):

want to see the options?

OpenStudy (debbieg):

if it's the inverse, switch and solve: change x for y and then solve for y. \[x=5y+12\]\[x-12=5y\]\[y=\frac{ x-12 }{5 }\]

OpenStudy (anonymous):

oh ok i get it thank you

OpenStudy (debbieg):

She was just asking how to evaluate f(-2) for a given function. I doubt she's doing derivatives?.....

OpenStudy (debbieg):

You're welcome. :)